OpenStudy (anonymous):

How far from home are you after 4 hours? Answer in units of m

OpenStudy (anonymous):

Driving toward home at a constant rate of 60 mph, you are 529 mi away. The distance, d, in miles away from home after t hours is given by d = 529 − 60

OpenStudy (anonymous):

How far from home are you after 4 hours? Answer in units of m

OpenStudy (anonymous):

check yr eqn... i think it shud be d = 529 - 60t

OpenStudy (anonymous):

If I am correct then your calculation shud be after 4 hours d = 529 - 60*4 = 529 - 240 = 289